A 13 Ohms coffee maker and a 16 Ohms frying pan are connected in series across a 120 V source of voltage. A 26 Ohms bread maker is also connected across the 120 V source and is in parallel with the series combination. Find the total current supplied by the source of voltage._______________A
1
Expert's answer
2012-02-14T09:46:22-0500
We need to use the rules of making equivalent resistors, and the Ohm's law.
The 14 and 16 ohm are in series, so let's add them up.
That sum, 30 ohm, is in parallel with the 23 ohm.
The equivalent resistance is the product divided by the sum, or 690 / 53 ohm, or about 13.02 ohms.

Now we can apply the Ohms law:

I = U/R = 120/13.02 = 9.21 A.